OpenStudy (anonymous):

Can I get help with this? Serena bought 3/4 pound of cheese. She used 2/3 of the cheese for sandwiches. How much cheese did Serena use for sandwiches?

OpenStudy (anonymous):

I believe it's 1/2 but I'm not sure

OpenStudy (anonymous):

3/4 = .75 2/3 of .75 = .50 I believe she would have used 1/2 as well.

OpenStudy (anonymous):

I'll try it as \[\frac{ 1 }{ 2 }\] Thanks for the reassurance @tbonds

OpenStudy (anonymous):

Glad I could help. :)

OpenStudy (anonymous):

It was 1/2 :D !